In a given rectangle,the length is 1 more than 2 times the width.the perimeter is 56.what are the dimensions ?
Answer by mananth(16946)

let width be x
length = 2x+1
..
Perimeter = 2*(L+W)
2*(x+2x+1)=56
2*(3x+1)=56
6x+2=56
add -2
6x+2-2=56-2
6x=54
/6
x= 9 the width
..
length = 2x+1
=2*9 +1
=19 the length
...
m.ananth@hotmail.ca